1.explain why it's not safe for passengers to stand on the upper decker in a double decker bus
2.a uniform rule is balanced at 30cm mark when a mass of 50g is hanging from it's zero mark.calculate the weight of the rule
1
Expert's answer
2020-04-06T08:44:28-0400

1)The topdeck standing passenger will sway outwards when the double-decker will turn.(On a Left turn, the standing passenger will sway towards the Right side of the bus)

This will cause the Centre of Gravity to move outwards (significantly more than the seated passenger) and away from the direction of the turn. It will be moving away from the track-width of the double-decker. This is the reason why no standing passengers are not allowed on the top deck of a double-decker bus.


2)

"W(0.5L-0.3L)=F(0.3L)"

"Mg(0.2)=mg(0.3)"

"M=1.5m=1.5(50)=150\ g"

"W=Mg=(0.15)(9.8)=1.5\ N"